OpenStudy (anonymous):

What is the solution of the system? Use substitution. 2x-y=-7 4x-y=-4

OpenStudy (ash2326):

@MissyK what's the coefficient of y in both equations?

OpenStudy (anonymous):

I don't know o_o It doesn't say..

OpenStudy (anonymous):

(–1.5, 4) (1.5, 10) (4, –1.5) (–1.5, –2)

OpenStudy (precal):

2x-y=-7 -y=-7-2x y=7+2x sub that into the second equation

OpenStudy (anonymous):

The answer choices.. sorry D:

OpenStudy (precal):

if you know the points just sub them into both equations, the one that comes out true for both is your solution

OpenStudy (anonymous):

I'm sorry.. I wouldn't be here if I knew that much.

OpenStudy (precal):